# DiRW: Path-Aware Digraph Learning for Heterophily

### v1

@@ A.7. Sensitivity Analysis
To systematically evaluate the influence of hyperparameters on the experimental outcomes, we performed an in-depth analysis focusing on two pivotal parameters: the minimum walk length and the minimum walk number . The findings of these experiments are depicted in Fig. 7, where all other parameters were kept constant to isolate the effects of and .
Our analysis reveals a notable trend of improvement in node classification performance as the increases. However, it is important to note that the increase in walk numbers, while beneficial for performance, cannot be pursued indefinitely due to the concomitant escalation in time and space complexity, which poses practical limitations. Regarding the minimal walk length, the impact on node classification performance is less straightforward. There is no definitive pattern of improvement or decline in node classification performance as is incremented. Interestingly, our results show that in homophilous digraphs, such as coraml and wikics, shorter walk lengths are more conducive to achieving better performance. In contrast, in heterophilous digraphs like Actor and Amazon-Rating, longer walk lengths are associated with superior performance. This dichotomy is in line with our prior discussions regarding the presence of high-order homophily in heterophilous graphs, which suggests that longer walks are essential for capturing the intricate relationships within these complex digraph structures.
A.8. Comparison of DiRW with Other PathGNNs
Substantial researches acknowledge the capability of path information in distilling the complex structural and homophilous patterns in graphs, thereby enhancing the characterization of node attributes in scenarios marked by heterophily. GeniePath (Liu et al., 2019) proposes an adaptive path layer which consists of two complementary functions designed for breadth and depth exploration. The breadth exploration learns the importance of different one-hop neighborhoods, while the depth exploration captures signals from neighbors of different hops away. SPAGAN (Yang et al., 2021) exploits the shortest path to accounts for the influence of a sequence of nodes between the center node and its higher-order neighbors. It also conducts path-based attention to obtain a more effective aggregation. PathNet (Sun et al., 2022) designs a maximal entropy-based path sampler to sample paths containing homophilous and structural context. It also introduces a structure-aware RNN to preserve order and distance information, learning the semantic contexts of neighborhoods. RAWGNN (Jin et al., 2022) utilizes Node2Vec (Grover and Leskovec, 2016) to simulate breadth first search and depth first search to capture homophily and heterophily information in graphs, respectively. It also introduces a new path-based aggregator based on RNN. PathMLP (Xie et al., 2023) observes high-order homophily embedded in heterophilous graphs and proposes a light-weight model based on MLP. It designs a similarity-based path sampling strategy to capture paths with high-order homophily for originating nodes.
Despite recent advancements, their walk strategies are tailored for undirected graphs and lack generalizability. Our empirical study in Fig. 1(a) and 1(b) have revealed that naive random walk leads to severe walk interruption problems, which indicates that the majority of walk sequences fail to gather extensive information. To address this challenge, DiRW designs a direction-aware walk strategy to identify potential relationships between the current node and its generalized neighbors through walk sequences. Furthermore, the complex relationships inherent in directed topology pose unique challenges to the naive random walks, necessitating further investigations to develop fine-grained walk rules. The empirical study in Fig. 2(a) and 2(b) demonstrates that the one-size-fits-all sampling approaches overlook the distinct contexts of nodes within complex topology. To solve this problem, DiRW fine-tunes the walk numbers and walk lengths on a node-adaptive level, providing a rich information source for representing the current node.
It is noteworthy that our proposed homophily entropy is fundamentally different from the maximal entropy random walk (MERW) mentioned in PathNet (Sun et al., 2022). The MERW makes use of the maximum entropy rate from information theory in (Cover, 1999), adjusting the walk probabilities to be proportional to the eigenvector centrality. On the other hand, the homophily entropy is designed to inform the walk length instead of guiding the walk probabilities. Furthermore, conceptually, the entropy in MERW depends on the transition probabilities of each node, which are solely determined by the graph’s topological structure. In contrast, the probability distribution used in homophily entropy is based on the features of nodes within a walk sequence, enabling it to measure the quality of the sequence.
A.9. PathGNN and Graph Attention
PathGNNs and graph attention mechanisms are two innovative approaches in graph representation learning, each with its unique methodology for capturing the nuances of graph-structured data.
Graph attention techniques focus on the local structure of graphs to allocate weights among nodes (Brody et al., 2022). The process typically unfolds in several distinct steps: (1) Computing node similarities. (2) Yielding attention weights. (3) Aggregating features with weights. The whole process effectively propagates and updates the node features in a manner that reflects their relative importance.
Similarly, PathGNNs systematically sample paths to capture the relationships between nodes. The operation of PathGNNs can be broken down into the following steps: (1) Sampling according to meticulously designed walk probabilities. (2) Forming path features by aggregating the node features along paths. (3) Deriving the node representations by integrating distinct path features. Broadly speaking, PathGNNs can be regarded as a lightweight variant of graph attention mechanisms. Unlike traditional attention models that may involve complex learning processes, PathGNNs rely on a more streamlined sampling procedure, which enhances their scalability. By avoiding the computational overhead associated with learning attention weights, PathGNNs offer a more flexible alternative for graph learning, particularly in heterophilous scenarios.

### v3

@@ 4.5. Sensitivity Analysis
To evaluate the impact of hyperparameters on performance, we conducted an in-depth analysis of the minimum walk length and the minimum walk number in Fig. 6. All other parameters were kept constant to isolate the effects of and .
Our analysis reveals a clear trend of improved performance as the minimum walk number increases. However, while more walks enhance performance, this cannot be pursued indefinitely due to the accompanying rise in time and space complexity.
The impact of is less straightforward. There is no obvious performance improvement or decrease as increases. Notably, in homophilous digraphs (e.g., Coraml, WikiCS), shorter walk lengths yield better performance. Conversely, in heterophilous digraphs (e.g., Actor, Rating), longer walk lengths are linked to superior performance. This aligns with our earlier discussions on high-order homophily in heterophilous graphs, indicating that longer walks are crucial for capturing complex relationships within these structures.
4.6. Path Quality Analysis
We conducted an in-depth analysis comparing the path quality sampled by our optimized path sampler with those generated by the DFS and BFS utilized in RAWGNN, measured by the homophily entropy in Eq. (7). The experimental outcome is shown in Fig. 7.
The results demonstrate that the paths sampled by DiRW consistently exhibit lower homophily entropy compared to those produced by DFS and BFS. This superiority is primarily attributed to the integrated consideration of both topology and feature in the walk probability, coupled with the adaptive walk length guided by homophily entropy. Furthermore, the DFS demonstrated the highest homophily entropy in most scenarios because BFS explores the immediate neighborhood, which aligns well with the homophily assumption. However, in scenarios characterized by strong heterophily (e.g., Actor), DFS outperforms BFS. DiRW achieves a harmonious balance between these two approaches, leveraging the strengths of both while mitigating their respective weaknesses.
5. Conclusion and Future Work
In this work, we emphasize the necessity of DiGNNs for modeling digraph-structured data and introduce the DiPathGNN mechanism to capture inherent homophilous information in digraphs. Our empirical analysis reveals the limitations of current PathGNNs, particularly their overlook of edge direction and coarse-grained sampling strategies. To overcome these challenges, we propose DiRW, a novel path-based digraph learning method characterized by its direction-aware path sampler, which is fine-tuned based on walk probability, length, and number. Additionally, DiRW utilizes a node-wise learnable path aggregator to create nuanced node representations. Experiments show that DiRW achieves state-of-the-art performance in node- and link-level tasks, particularly in heterophilous scenarios, providing an efficient and robust solution for digraph learning. Looking ahead, several promising directions for future research emerge. First, we plan to explore more efficient sampling strategies. Then, developing adaptive sampling methods would also improve model performance. Lastly, designing improved learning mechanisms to capture the complex relationships in digraphs may yield more accurate predictions. These future efforts will deepen our understanding of digraph learning and could lead to breakthroughs in modeling complex digraph-structured data.
Acknowledgements.
This work is supported by the NSFC Grants U2241211, 62427808, and U24A20255. Rong-Hua Li is the corresponding author.GenAI Usage Disclosure
We used ChatGPT exclusively for language polishing in this work. We take full responsibility for the integrity of the work’s intellectual content and its adherence to academic ethical standards.